OpenStudy (anonymous):

If the populATION of bacteria doubles every 30 minutes, how long would it take for the population to triple? (question from logarithms homework)

OpenStudy (jwt625):

x=0,y=1;x=30,y=2,so \[y=(\sqrt[30]{2})^x\] when y=3,solve for x:\[3=(\sqrt[30]{2})^x\] \[x=\log_{\sqrt[30]{2}}3 \]

OpenStudy (anonymous):

ok or try \[3=2^{\frac{t}{30}}\] solve for t

OpenStudy (jwt625):

\[x=30*\log_{2} 3\]

OpenStudy (anonymous):

solve via \[\ln(3)=\frac{t}{30}\ln(2)\] \[\frac{t}{30}=\frac{\ln(3)}{\ln(2)}\] \[t=30\frac{\ln(3)}{\ln(2)}\]

OpenStudy (anonymous):

same answer as jwt but this one you can find the number on a calculator.

OpenStudy (anonymous):

i get 47.55 rounded so between day 47 and 48

OpenStudy (jwt625):

between 47and48 minutes